a ball is projected with initial velocity of 18 m/s at an angle of 25 degree. what is the time of flight of the ball?


1
Expert's answer
2021-11-25T08:37:23-0500

The time of flight is given as follows (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"T = \dfrac{2u\sin\theta }{g}"

where "u=18m/s, \theta =25\degree" and "g = 9.8m\/s^2" is the gravitational acceleration. Thus, obtain:


"T = \dfrac{2\cdot 18m/s\cdot \sin 25\degree}{9.8m/s^2} \approx 1.6s"

Answer. 1.6s.